OpenStudy (anonymous):

what is the d^2y/dx^2 of dy/dx = -x^3/(16-x^4)^(3/4)

OpenStudy (anonymous):

If your trying to find the second derivaitve, i believe you should use the quotient rule again

OpenStudy (anonymous):

maybe it would be easier to start with \[y'=-\frac{x^3}{y^3}\] to find the second derivative

OpenStudy (anonymous):

again it depends on whether you care if you have a y in your answer. you pick. we could also use the quotient rule here \[y'=\frac{x^3}{\sqrt[4]{(16-x^4)^3}}\]

OpenStudy (anonymous):

notice that these are identical because the denominator of the second one is just y^3

OpenStudy (anonymous):

you still need an answer?